OpenStudy (anonymous):

The height of an object dropped from a hot-air balloon can be determined by the formula h=-16t^2+s, where ( h ) is the height of the object in feet, ( t ) is the elapsed time in seconds, and ( s ) is the height of the balloon. If the balloon is 400 feet from the ground, how long will it take for an object dropped from the balloon to hit the ground? a) 2.5 seconds b) 5 seconds c) 7.5 seconds d) 10 seconds

OpenStudy (anonymous):

If the balloon is 400 feet up. Replace s with 400 giving h=-16t^2+400. Since you want the time when it hits the ground the height of the ball(h) is ZERO. So: 0=-16t^2+400. Solve for t. (Be weary, you will get a plus/minus from the squared, disregard the negative one. Negative time makes no sense :P)

OpenStudy (anonymous):

And I got 5. :P

OpenStudy (anonymous):

i got B but im not sure. I used the formula and plugged in h= -16(5)^2 + 400 and brought it down to h= -16(25) + 400 h= -400+400 h=0 so im caught between B and D

OpenStudy (anonymous):

Exactly. That means, the height is zero relative to the GROUND. Which is what it wants :P